Acute gastrointestinal illness was reported to school health department soon after eating contaminated vanilla ice cream at school picnic. Calculate the attack rate, given that; of those who ate the food, 43 were ill and 11 were not ill. Of those who did not eat the food 3 were ill and 18 were not ill. Calculate the ratio of the attack rate between those ate and those didn't and explain their indication/s
Answer #1

Attack rate = total number of people who are I'll decided by total no of people who are at risk

According to give question attack rate of vanilla eating is =43/54=.79

Attack rate of not eat is =3/21=.14

Ratio of attack rate is .70/.14 =5.64